OpenStudy (faiqraees):

√(1+x)*√(1-x) Can it be simplified?

OpenStudy (faiqraees):

@ganeshie8

OpenStudy (anonymous):

$\sqrt{1+x}\left(\sqrt{1-x}\right)=\sqrt{x+1}\sqrt{-x+1}$√1+x(√1−x)=√x+1√−x+1 Steps $\sqrt{1+x}\left(\sqrt{1-x}\right)$√1+x(√1−x) $\mathrm{Remove\:parentheses}:\quad\left(a\right)=a$Remove parentheses: (a)=a $=\sqrt{1+x}\sqrt{1-x}$=√1+x√1−x

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

its no

OpenStudy (faiqraees):

@usercode3rror Can you use the equation editor?

OpenStudy (anonymous):

@faiqraees where would i find this